OpenStudy (anonymous):

the resistannce of a machine winding at 15 degrees clecious is 150 ohms and at 76.6 degrees celcious is 185 ohms. determine the reduction in the coil current due to the temperature rise, assuming that the winding is fed from a constant d.c supply

OpenStudy (radar):

I=E/R or\[I=ER ^{-1}\] \[dI/dR = -ER ^{-2}=-E/R ^{2}\]